Use synthetic division to determine whether the numbers are zeros of the polynomial function.
-3,1; h(x)=x^4+3x^3+2x^2-4x-2
is -3 a zero of this function?
is 1 a zero of this function?

Answer by Alan3354(69443)

Use synthetic division to determine whether the numbers are zeros of the polynomial function.
-3,1; h(x)=x^4+3x^3+2x^2-4x-2
-------------------
Just sub -3 for x, see if it's a zero
h(-3) = 81 - 81 + 18 + 12 - 2 = 28, not a zero
-----------
h(1) = 1 + 3 + 2 - 4 - 2 = 0